OpenStudy (anonymous):

Simplify the expression 5^3 × 5^8.

OpenStudy (solomonzelman):

\(\Large\color{black}{ 5^3\times 5^8=5^{3+8}=5^{11}}\)

OpenStudy (solomonzelman):

=48828125

OpenStudy (solomonzelman):

You don't need the last part, just 5^11.